OpenStudy (anonymous):

a problem for u Two equations in a plane are given a line y=2b(a-x) a parabola y=x^2-2ax+1

OpenStudy (anonymous):

That isn't a question or a problem.

OpenStudy (anonymous):

\[A=\left\{ (a,b) \in R^2 \ such \ that \ \ line \ and \ Parabola \ dont \ intercept \ \right\}\]

OpenStudy (anonymous):

find the area of A

OpenStudy (anonymous):

intersect*

OpenStudy (anonymous):

I don't completely understand.

OpenStudy (anonymous):

A is a set of points (a,b) such that line and parabola dont intersect each other problem is finding area of A

OpenStudy (anonymous):

It's the area of the part where they don't intersect?

OpenStudy (anonymous):

can you draw it out? And 'show' what you mean.

OpenStudy (anonymous):

i cant draw it because a and b are unknown

OpenStudy (anonymous):

solution would be 2ab < 1...this are the pairs of a and b

OpenStudy (anonymous):

answer is a^2+b^2<=1

OpenStudy (anonymous):

@arghya \[2b(a-x)=x^2-2ax+1 \\ x^2+2(b-a)x-2ab+1=0 \]

OpenStudy (anonymous):

Yeah then ya will take the discrimant less than zero..

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

I did the same..but did a silly mistake,so was not getting

OpenStudy (anonymous):

\[\Delta<0 \\ (2(b-a))^2-4(-2ab+1)<0 \\ 4(a^2+b^2-4ab)+8ab-4<0 \\ 4(a^2+b^2)<4 \\ a^2+b^2<1 \\ then \ the \ area \ of \ A \ is \ \frac{\pi}{4}\]